# A Comprehensive Survey of Website Fingerprinting Attacks and Defenses in Tor: Advances and Open Challenges

### v1

@@ V-B4 Adversarial Perturbation-Based defenses
Dolos [172]
KNN [9] NA >>98.00
@@
k-FP [36] NA >>98.00
CUMUL [11] NA >>97.00
DF [93] NA >>96.00
Var-CNN [96] NA >>95.00
@@
In recent years, DL-based WF attacks represent a significant evolution in the ability to de-anonymize users based on their encrypted network traffic. By automating the feature extraction process and leveraging powerful models like CNNs, RNNs, and LSTMs, these attacks can achieve very high accuracy in identifying websites, even under challenging conditions like traffic obfuscation [175, 176]. Over the years, adversarial perturbations, derived from the field of adversarial ML/DL, refer to carefully designed modifications of input data that cause ML/DL models to make incorrect predictions [177, 178]. Figure 10 presents an overview of the adversarial perturbation-based mechanism. The adversarial perturbation–based mechanism adds minimal, often imperceptible, noise or modifications to a legitimate input and thereby forces a target classifier to produce an incorrect prediction. This subsection will explore adversarial perturbation-based defenses for WF, comparing their methods, effectiveness, strengths, and drawbacks. The comparison of the representative adversarial perturbation-based defenses is summarized in TABLE XIII. TABLE XIII presents the effectiveness of defenses, showing changes in WF attack accuracy (attack rate) and defense success rate (defense rate). The strengths and drawbacks of each defense model are also summarized in TABLE XIII.
@@
Previous defenses like traffic morphing [45] and adaptive padding [58, 69] are limited by static patterns and high detectability. To solve this issue, Imani et al. [179] presented an Adversarial Traces defense method by adding padding to a Tor traffic trace in a manner that reliably fools the classifier into classifying it as coming from a different site. Li et al. [108] addressed the challenge of evading internet censorship through traffic analysis attacks, such as WF and protocol fingerprinting (PF). The authors propose FlowGAN, which uses Generative Adversarial Networks (GANs) to dynamically camouflage censored traffic as benign network flows, thereby bypassing censorship. Cai et al. [180] introduced JumpEstimate, a black-box countermeasure against WF attacks, which leverages adversarial machine learning to confuse the decision boundaries of DL and ML-based WF classifiers. Without relying on manually extracted features, JumpEstimate uses adversarial traffic generated through an improved boundary attack and Monte Carlo estimation to maximize confusion while resisting retraining by the attacker.
@@
Deep neural networks are known to be vulnerable to adversarial examples. Hou et al. [59] introduced WF-GAN, which utilizes adversarial learning techniques to generate adversarial examples for WF classifiers. The authors demonstrate that WF-GAN effectively generates adversarial examples to obscure traffic patterns, achieving a 90% success rate with minimal overhead (5-15%) in both untargeted and targeted defense scenarios. The results indicate that WF-GAN outperforms existing defenses, such as Walkie-Talkie, providing a robust solution for protecting encrypted traffic against sophisticated WF attacks. Rahman et al. [60] introduced Mockingbird, a technique designed to generate traces that resist adversarial training by introducing randomness into the process. Instead of following predictable gradients, Mockingbird explores the space of viable traces through random movements, making it more challenging for adversarially trained classifiers to defend against. Unlike conventional adversarial techniques, which are typically vulnerable to adversarial retraining, Mockingbird evades such defenses by perturbing network traces in a way that progressively shifts the source trace toward multiple randomly selected target traces. This approach does not rely on gradient-based optimization, complicating adversarial adaptation and defense. To address the limitation of requiring the entire original traffic trace for adversarial perturbations in WF defenses, Jiang et al. [170] introduced PST, a practical adversarial learning-based defense against WF attacks, which dynamically generates perturbations using only the initial bursts of network traffic rather than requiring the entire trace. PST employs an encoder-decoder neural network with attention mechanisms to predict subsequent traffic bursts, then applies a modified Boundary Attack to compute adversarial perturbations, and transfers these perturbations to the remaining traffic. Evaluation results show that PST effectively disrupts the network traffic pattern, achieving a high evasion rate of 87.6%. This outperforms Walkie-Talkie by over 31.59% at the same bandwidth overhead, while only requiring observation of 10 transferred bursts.
@@
Later, Nasr et al. [61] introduced an adversarial approach to counter deep neural network (DNN)-based traffic analysis techniques. The key idea behind their adversarial perturbations algorithm is to generate “blind” perturbations, which are independent of the target inputs, by solving specific optimization problems. This method operates in real-time, without requiring prior knowledge of network traffic, making it highly practical. It perturbs live network traffic on-the-fly, without the need to buffer packets or rely on any prior understanding of traffic patterns. Building on Nasr’s work, Shan et al. [172] proposed Dolos, a defense against WF attacks that can be applied in real-time to network traffic. Dolos uses parameterized randomness to thwart adaptive attacks, including adversarial training. A key element of Dolos is the use of trace-agnostic adversarial patches, generated via a user-side secret, to protect network traces. Unlike adversarial examples tailored for specific inputs, these patches are pre-computed sequences of dummy packets that can induce misclassifications across a broad range of inputs, effectively safeguarding visits to specific websites. Hou et al. [20] proposed a universal adversarial defense against WF attacks by generating perturbations that deceive DNN-based classifiers, without requiring prior knowledge of the full traffic trace. The method utilizes the Jacobian matrix to estimate the effect of dummy packet insertions on classifier outputs, facilitating efficient perturbation selection. In the same year, Hou et al. [181] introduced another defense mechanism, Attack to Attack (A2A), which uses adversarial examples to disrupt the attacker’s classifier. A2A operates as a proxy near the client, perturbing encrypted traffic by iteratively inserting dummy packets based on feedback from a substitute model trained to mimic the attacker’s classifier. Evaluated on a public Tor dataset [93, 94], A2A reduces the attacker’s accuracy to 57.6% with only 2.2% bandwidth overhead, significantly outperforming existing defenses like WTF-PAD and Walkie-Talkie, which incur higher overheads (31–64%). Wang et al. [182] introduced Local Optimal Mutation (LOM), a method that mutates sensitive traffic traces toward target populations of non-sensitive traces using a gradient-descent-inspired strategy. This approach minimizes computational costs while maintaining effectiveness. By combining the benefits of adversarial samples with a WF defense method based on local optimal mutation, LOM achieves the generation of adversarial traffic traces with shorter preparation times.
@@
Inspired by the universal adversarial Perturbations (UAPs) mechanism [177], Sun et al. [21] developed a GAN-based approach for UAP generation, called WF-UAP. Leveraging the power of GANs, this method efficiently creates UAPs that make adversarial traces harder to distinguish from original traffic. The crafted UAPs are independent of target network traffic traces and can disrupt the classifiers used in WF attacks when applied to any traffic trace. In addition, this method can inject dummy packets into live traffic, making adversarial traces indistinguishable from real ones while minimizing bandwidth overhead. Adversarial perturbations have excellent defense capabilities against WF and low bandwidth overhead, but they need to get the complete website traffic to generate defense data, which is obviously impractical. Tang et al. [183] proposed segmented adversary defense (SAD) for deep learning-based WF attacks. In SAD, sequence data are divided into multiple segments to ensure that SAD is feasible in real scenarios. The adversarial examples for each segment of data can be generated by SAD. Finally, dummy packets are inserted after each segment original data.
@@
Gong et al. [46] proposed a solution using Generative Adversarial Networks (GANs) to generate realistic traffic patterns that mimic the behavior of various web pages. These synthetic traces are used to regulate packet transmission, obscuring the website being visited by introducing randomness in both packet sizes and sending times. This dynamic adjustment of packet bursts based on buffered data makes it significantly harder for attackers to distinguish between different websites. Jiang et al. [171] introduced Targeted LightwEight Defense (TED), a lightweight and efficient method to defend against WF attacks in the Tor network. TED introduces trace-level defenses, customizing the defense for each unique trace—even for the same website. In addition, TED leverages a fully convolutional network (FCN) combined with adapted Grad-CAM to extract key discriminative features from traffic traces. These are the elements deep learning-based WF attacks rely on. TED then precisely perturbs these key features by injecting a small number of dummy packets, misleading the attacker’s classifier without heavy overhead.
@@
Many previous adversarial perturbation-based WF defenses have been criticized for relying on unrealistic assumptions, such as full access to the target model or needing to operate on the entire website trace. To overcome these limitations, Li et al. [74] introduced Minipatch that strategically injects few dummy packets in real-time traffic to evade the attacker’s classifier. Unlike previous adversarial-based defenses, which suffer from high bandwidth overhead or unrealistic assumptions (e.g., requiring white-box access to the attack model), Minipatch operates in a black-box setting and maintains low overhead (less than 5%). Bai et al. [184] introduced WGAN-GA that combines Wasserstein Generative Adversarial Networks (WGAN) with Genetic Algorithms (GA) to generate adversarial traffic samples that obfuscate fingerprinting attempts. Holland et al. [185] employed a Generative Adversarial Network (GAN)-inspired architecture to ensure the defense adapts to evolving attacker strategies. Yin et al. [186] introduced Greedy Injection Attack (GIA) and Sniper, two adversarial example-based WFD frameworks. These methods dynamically perturb network traffic to mislead DL-based WFAs with minimal bandwidth overhead. Prior WF defenses often fail to enforce real-time constraints or require symmetric proxy deployments. To address this issue, Li et al. [187] exploited patterns in packet timing and size to compromise user privacy. Then, the authors proposed Prism, a lightweight, asymmetric framework that dynamically perturbs live traffic to evade detection while maintaining transmission integrity. Qiao et al. [173] proposed a black-box WF defense called Acup3, which operates under the assumption that the attacker’s classifier is inaccessible and unknown. This technique makes traffic traces from different websites appear more similar, increasing the difficulty of classification. Acup3 generates perturbations without needing access to the actual traffic traces, making it practical for real-world deployment. By diversifying the perturbations applied to traffic traces of different users visiting the same website, Acup3 reduces the effectiveness of AT, as the classifier cannot easily generalize from previously seen perturbations.
@@
To mitigate the high bandwidth and latency overhead of prior WF defenses, Jiang et al. [174] introduced KimeraPad that uses deep reinforcement learning (DRL) to generate adversarial perturbations in real-time, making it difficult for attackers to classify traffic patterns accurately. KimeraPAD can be implemented in real-time without requiring prior knowledge of the exact packet sequences or their lengths, making it practical for real-world use. Wang et al. [188] proposed CMAES-WFD, a black-box adversarial defense against WF attacks, which leverages the Covariance Matrix Adaptation Evolution Strategy (CMAES) to generate perturbations that mislead deep neural network (DNN)-based classifiers while minimizing bandwidth overhead. The method transforms perturbation generation into a constrained optimization problem, injecting dummy packets into critical positions of traffic bursts to disrupt classification without requiring access to the target model’s internal parameters.
@@
Existing defenses, particularly those based on Adversarial Examples (AEs) [60, 189], face challenges such as requiring complete traffic traces for perturbation calculation and being vulnerable to Adversarial Training (AT) by attackers. To address this issue, Qiao et al. [47] presented a trace-Agnostic LightwEight Transferable Robust (ALERT) WF defense to resist DNN-based WF attacks. ALERT uses a GAN-based generator to create universal perturbations without requiring access to real-time traffic traces, enabling packet manipulation before browsing starts. It incorporates a random website imitation strategy, where each user’s traffic is perturbed to mimic a randomly selected website, introducing high diversity and mitigating adversarial training (AT) effectiveness. In closed-world experiments, ALERT reduces attack success rates (ASRs) below 13% for strong WF attacks like DF. In open-world scenarios, ALERT achieves the best trade-off between true and false positive rates while maintaining minimal communication overhead ( 19.9%). Jiang et al. [48] introduced RUDOLF, an efficient and adaptive WF defense based on Soft Actor-Critic (SAC) reinforcement learning. RUDOLF dynamically adds dummy packets in real-time by training an RL agent to make burst-level decisions, eliminating the need for complete traffic traces and making it practical and adaptive. The agent receives feedback via a dual-component reward function that balances evasion success against bandwidth overhead, allowing it to learn optimal perturbation strategies. Evaluations show RUDOLF achieves attack accuracies as low as 15–20% with less than 30% bandwidth overhead.

### v2

@@ V-B4 Adversarial Perturbation-Based defenses
Dolos [shan2021patch]
KNN [wang2014effective] NA >>98.00
@@
k-FP [hayes2016k] NA >>98.00
CUMUL [panchenko2016website] NA >>97.00
DF [sirinam2018deep] NA >>96.00
Var-CNN [bhat2018var] NA >>95.00
@@
In recent years, DL-based WF attacks represent a significant evolution in the ability to de-anonymize users based on their encrypted network traffic. By automating the feature extraction process and leveraging powerful models like CNNs, RNNs, and LSTMs, these attacks can achieve very high accuracy in identifying websites, even under challenging conditions like traffic obfuscation [holland2024tor, liu2022advtraffic]. Over the years, adversarial perturbations, derived from the field of adversarial ML/DL, refer to carefully designed modifications of input data that cause ML/DL models to make incorrect predictions [moosavi2017universal, huang2025wf]. Figure 10 presents an overview of the adversarial perturbation-based mechanism. The adversarial perturbation–based mechanism adds minimal, often imperceptible, noise or modifications to a legitimate input and thereby forces a target classifier to produce an incorrect prediction. This subsection will explore adversarial perturbation-based defenses for WF, comparing their methods, effectiveness, strengths, and drawbacks. The comparison of the representative adversarial perturbation-based defenses is summarized in TABLE XIII. TABLE XIII presents the effectiveness of defenses, showing changes in WF attack accuracy (attack rate) and defense success rate (defense rate). The strengths and drawbacks of each defense model are also summarized in TABLE XIII.
@@
Previous defenses like traffic morphing [wang2017walkie] and adaptive padding [dyer2012peek, cai2014cs] are limited by static patterns and high detectability. To solve this issue, Imani et al. [imani2018adversarial] presented an Adversarial Traces defense method by adding padding to a Tor traffic trace in a manner that reliably fools the classifier into classifying it as coming from a different site. Li et al. [li2019dynamic] addressed the challenge of evading internet censorship through traffic analysis attacks, such as WF and protocol fingerprinting (PF). The authors propose FlowGAN, which uses Generative Adversarial Networks (GANs) to dynamically camouflage censored traffic as benign network flows, thereby bypassing censorship. Cai et al. [cai2020jumpestimate] introduced JumpEstimate, a black-box countermeasure against WF attacks, which leverages adversarial machine learning to confuse the decision boundaries of DL and ML-based WF classifiers. Without relying on manually extracted features, JumpEstimate uses adversarial traffic generated through an improved boundary attack and Monte Carlo estimation to maximize confusion while resisting retraining by the attacker.
@@
Deep neural networks are known to be vulnerable to adversarial examples. Hou et al. [hou2020wf] introduced WF-GAN, which utilizes adversarial learning techniques to generate adversarial examples for WF classifiers. The authors demonstrate that WF-GAN effectively generates adversarial examples to obscure traffic patterns, achieving a 90% success rate with minimal overhead (5-15%) in both untargeted and targeted defense scenarios. The results indicate that WF-GAN outperforms existing defenses, such as Walkie-Talkie, providing a robust solution for protecting encrypted traffic against sophisticated WF attacks. Rahman et al. [rahman2020mockingbird] introduced Mockingbird, a technique designed to generate traces that resist adversarial training by introducing randomness into the process. Instead of following predictable gradients, Mockingbird explores the space of viable traces through random movements, making it more challenging for adversarially trained classifiers to defend against. Unlike conventional adversarial techniques, which are typically vulnerable to adversarial retraining, Mockingbird evades such defenses by perturbing network traces in a way that progressively shifts the source trace toward multiple randomly selected target traces. This approach does not rely on gradient-based optimization, complicating adversarial adaptation and defense. To address the limitation of requiring the entire original traffic trace for adversarial perturbations in WF defenses, Jiang et al. [jiang2020pst] introduced PST, a practical adversarial learning-based defense against WF attacks, which dynamically generates perturbations using only the initial bursts of network traffic rather than requiring the entire trace. PST employs an encoder-decoder neural network with attention mechanisms to predict subsequent traffic bursts, then applies a modified Boundary Attack to compute adversarial perturbations, and transfers these perturbations to the remaining traffic. Evaluation results show that PST effectively disrupts the network traffic pattern, achieving a high evasion rate of 87.6%. This outperforms Walkie-Talkie by over 31.59% at the same bandwidth overhead, while only requiring observation of 10 transferred bursts.
@@
Later, Nasr et al. [nasr2021defeating] introduced an adversarial approach to counter deep neural network (DNN)-based traffic analysis techniques. The key idea behind their adversarial perturbations algorithm is to generate “blind” perturbations, which are independent of the target inputs, by solving specific optimization problems. This method operates in real-time, without requiring prior knowledge of network traffic, making it highly practical. It perturbs live network traffic on-the-fly, without the need to buffer packets or rely on any prior understanding of traffic patterns. Building on Nasr’s work, Shan et al. [shan2021patch] proposed Dolos, a defense against WF attacks that can be applied in real-time to network traffic. Dolos uses parameterized randomness to thwart adaptive attacks, including adversarial training. A key element of Dolos is the use of trace-agnostic adversarial patches, generated via a user-side secret, to protect network traces. Unlike adversarial examples tailored for specific inputs, these patches are pre-computed sequences of dummy packets that can induce misclassifications across a broad range of inputs, effectively safeguarding visits to specific websites. Hou et al. [hou2021universal] proposed a universal adversarial defense against WF attacks by generating perturbations that deceive DNN-based classifiers, without requiring prior knowledge of the full traffic trace. The method utilizes the Jacobian matrix to estimate the effect of dummy packet insertions on classifier outputs, facilitating efficient perturbation selection. In the same year, Hou et al. [hou2021attack] introduced another defense mechanism, Attack to Attack (A2A), which uses adversarial examples to disrupt the attacker’s classifier. A2A operates as a proxy near the client, perturbing encrypted traffic by iteratively inserting dummy packets based on feedback from a substitute model trained to mimic the attacker’s classifier. Evaluated on a public Tor dataset [sirinam2018deep, rimmer2017automated], A2A reduces the attacker’s accuracy to 57.6% with only 2.2% bandwidth overhead, significantly outperforming existing defenses like WTF-PAD and Walkie-Talkie, which incur higher overheads (31–64%). Wang et al. [wang2022lom] introduced Local Optimal Mutation (LOM), a method that mutates sensitive traffic traces toward target populations of non-sensitive traces using a gradient-descent-inspired strategy. This approach minimizes computational costs while maintaining effectiveness. By combining the benefits of adversarial samples with a WF defense method based on local optimal mutation, LOM achieves the generation of adversarial traffic traces with shorter preparation times.
@@
Inspired by the universal adversarial Perturbations (UAPs) mechanism [moosavi2017universal], Sun et al. [sun2022practical] developed a GAN-based approach for UAP generation, called WF-UAP. Leveraging the power of GANs, this method efficiently creates UAPs that make adversarial traces harder to distinguish from original traffic. The crafted UAPs are independent of target network traffic traces and can disrupt the classifiers used in WF attacks when applied to any traffic trace. In addition, this method can inject dummy packets into live traffic, making adversarial traces indistinguishable from real ones while minimizing bandwidth overhead. Adversarial perturbations have excellent defense capabilities against WF and low bandwidth overhead, but they need to get the complete website traffic to generate defense data, which is obviously impractical. Tang et al. [tang2022sad] proposed segmented adversary defense (SAD) for deep learning-based WF attacks. In SAD, sequence data are divided into multiple segments to ensure that SAD is feasible in real scenarios. The adversarial examples for each segment of data can be generated by SAD. Finally, dummy packets are inserted after each segment original data.
@@
Gong et al. [gong2022surakav] proposed a solution using Generative Adversarial Networks (GANs) to generate realistic traffic patterns that mimic the behavior of various web pages. These synthetic traces are used to regulate packet transmission, obscuring the website being visited by introducing randomness in both packet sizes and sending times. This dynamic adjustment of packet bursts based on buffered data makes it significantly harder for attackers to distinguish between different websites. Jiang et al. [jiang2022effective] introduced Targeted LightwEight Defense (TED), a lightweight and efficient method to defend against WF attacks in the Tor network. TED introduces trace-level defenses, customizing the defense for each unique trace—even for the same website. In addition, TED leverages a fully convolutional network (FCN) combined with adapted Grad-CAM to extract key discriminative features from traffic traces. These are the elements deep learning-based WF attacks rely on. TED then precisely perturbs these key features by injecting a small number of dummy packets, misleading the attacker’s classifier without heavy overhead.
@@
Many previous adversarial perturbation-based WF defenses have been criticized for relying on unrealistic assumptions, such as full access to the target model or needing to operate on the entire website trace. To overcome these limitations, Li et al. [li2022minipatch] introduced Minipatch that strategically injects few dummy packets in real-time traffic to evade the attacker’s classifier. Unlike previous adversarial-based defenses, which suffer from high bandwidth overhead or unrealistic assumptions (e.g., requiring white-box access to the attack model), Minipatch operates in a black-box setting and maintains low overhead (less than 5%). Bai et al. [bai2023web] introduced WGAN-GA that combines Wasserstein Generative Adversarial Networks (WGAN) with Genetic Algorithms (GA) to generate adversarial traffic samples that obfuscate fingerprinting attempts. Holland et al. [holland2023detorrent] employed a Generative Adversarial Network (GAN)-inspired architecture to ensure the defense adapts to evolving attacker strategies. Yin et al. [yin2023defeating] introduced Greedy Injection Attack (GIA) and Sniper, two adversarial example-based WFD frameworks. These methods dynamically perturb network traffic to mislead DL-based WFAs with minimal bandwidth overhead. Prior WF defenses often fail to enforce real-time constraints or require symmetric proxy deployments. To address this issue, Li et al. [li2023prism] exploited patterns in packet timing and size to compromise user privacy. Then, the authors proposed Prism, a lightweight, asymmetric framework that dynamically perturbs live traffic to evade detection while maintaining transmission integrity. Qiao et al. [qiao2023resisting] proposed a black-box WF defense called Acup3, which operates under the assumption that the attacker’s classifier is inaccessible and unknown. This technique makes traffic traces from different websites appear more similar, increasing the difficulty of classification. Acup3 generates perturbations without needing access to the actual traffic traces, making it practical for real-world deployment. By diversifying the perturbations applied to traffic traces of different users visiting the same website, Acup3 reduces the effectiveness of AT, as the classifier cannot easily generalize from previously seen perturbations.
@@
To mitigate the high bandwidth and latency overhead of prior WF defenses, Jiang et al. [jiang2024kimerapad] introduced KimeraPad that uses deep reinforcement learning (DRL) to generate adversarial perturbations in real-time, making it difficult for attackers to classify traffic patterns accurately. KimeraPAD can be implemented in real-time without requiring prior knowledge of the exact packet sequences or their lengths, making it practical for real-world use. Wang et al. [wang2024cmaes] proposed CMAES-WFD, a black-box adversarial defense against WF attacks, which leverages the Covariance Matrix Adaptation Evolution Strategy (CMAES) to generate perturbations that mislead deep neural network (DNN)-based classifiers while minimizing bandwidth overhead. The method transforms perturbation generation into a constrained optimization problem, injecting dummy packets into critical positions of traffic bursts to disrupt classification without requiring access to the target model’s internal parameters.
@@
Existing defenses, particularly those based on Adversarial Examples (AEs) [rahman2020mockingbird, sadeghzadeh2021awa], face challenges such as requiring complete traffic traces for perturbation calculation and being vulnerable to Adversarial Training (AT) by attackers. To address this issue, Qiao et al. [qiao2024trace] presented a trace-Agnostic LightwEight Transferable Robust (ALERT) WF defense to resist DNN-based WF attacks. ALERT uses a GAN-based generator to create universal perturbations without requiring access to real-time traffic traces, enabling packet manipulation before browsing starts. It incorporates a random website imitation strategy, where each user’s traffic is perturbed to mimic a randomly selected website, introducing high diversity and mitigating adversarial training (AT) effectiveness. In closed-world experiments, ALERT reduces attack success rates (ASRs) below 13% for strong WF attacks like DF. In open-world scenarios, ALERT achieves the best trade-off between true and false positive rates while maintaining minimal communication overhead ( 19.9%). Jiang et al. [jiang2024rudolf] introduced RUDOLF, an efficient and adaptive WF defense based on Soft Actor-Critic (SAC) reinforcement learning. RUDOLF dynamically adds dummy packets in real-time by training an RL agent to make burst-level decisions, eliminating the need for complete traffic traces and making it practical and adaptive. The agent receives feedback via a dual-component reward function that balances evasion success against bandwidth overhead, allowing it to learn optimal perturbation strategies. Evaluations show RUDOLF achieves attack accuracies as low as 15–20% with less than 30% bandwidth overhead.